Two cards are drawn with replacement from an ordinary deck of 52 playing cards. What is the probability that both cards are of the same value (for example, both cards are twos or queens or aces)?
Answer by stanbon(75887)

Two cards are drawn with replacement from an ordinary deck of 52 playing cards. What is the probability that both cards are of the same value (for example, both cards are twos or queens or aces)?
---
How to succeed:
Pick a card type: 13 ways
Pick 2 cards of that type: 4C2 = 6
----------------------------------------
# of ways to succeed: 13*6 = 78
-----
# of possible selections: 52C2 = 650
------
So, P(both cards are same value) = 78/650 = 0.12
===================================================
Cheers,
Stan H.